OpenStudy (anonymous):

Two angles are complimentary. The sum of the measure of the first angle and half the second angle is 84. What are the measures of the first and second angles?

OpenStudy (anonymous):

Let the first angle be x and the second angle be y. Then x + y = 90 x + (1/2)y = 84 Solve these simultaneous equations to get the answers.

OpenStudy (anonymous):

I am so terrible with math. Would the first angle be 90 and the second 168

OpenStudy (anonymous):

84/2 is 42. 42/2 is 21. 42+21= 63. 84-63 = 21. The measure of each angle is 63 degrees.

OpenStudy (radar):

Stay with sklee's equations. The problem is set up properly. Proceed now with the solution. We have two equations, two unknowns. Multiply the (x+1/2 y = 84) by 2 getting: 2x+y = 168 now subtrtact this equation from his firs equation: x + y = 90 2x+y = 168 -x =-78 or x =78, the other angle is 90-78 = 12. The two angles are 12 and 78 degrees. They are complimentary, adding up 90 degrees and the 12 degree when diided by 2 and added to the 78 degree angle equals 84 degrees. It checks out.

OpenStudy (radar):

A good businessw would always check out and verify!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

i support Radar